OpenStudy (anonymous):

solve the system x-y=33x -8y=9 what's the value of y? Please help ! (: thank you A.-2 B.6 C.0 D.3

OpenStudy (anonymous):

Is this what the equations should look like? x-y=33x -8y=9

OpenStudy (anonymous):

If so, I would solve the top equation for y and substitute.

OpenStudy (anonymous):

I'm pretty sure its all in one that is what my page says but I will try that I still don't understand though!

OpenStudy (anonymous):

It can't be all one equation since you have 2 = signs. Do you know how to solve an equation for y?

OpenStudy (anonymous):

okay and no I don't

OpenStudy (anonymous):

I suggest solving the top equation because y has a coefficient of 1 and therefore will not result in a fraction.

OpenStudy (anonymous):

Use inverse operations to isolate y on one side of the =

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

To solve the following equation for y: 6x + 2y = 12 Isolate the y term by subtracting the x (2x) term from both sides. 2y = -6x + 12 Then divide both sides by 2: y= -3x + 6

OpenStudy (anonymous):

So is it -2

OpenStudy (anonymous):

It's hard to say. The original equations you gave me are not correct. If you can retype the equations, I can check your answer.

OpenStudy (anonymous):

x-y=33 x-8y=9

OpenStudy (anonymous):

That makes more sense. One moment please...

OpenStudy (anonymous):

what is the value of y I thought it was -2

OpenStudy (texaschic101):

that made a difference

OpenStudy (texaschic101):

x - y = 33 --> x = 33 + y now sub 33 + y in for x in the other equation and solve for y

OpenStudy (anonymous):

i am going to bet it is \[x-y=3\\3x -8y=9 \]

OpenStudy (anonymous):

just a guess mind you, i could be wrong, but that is what i think it might be

OpenStudy (anonymous):

You may be right.

OpenStudy (texaschic101):

check your problem again Kimmeeboo...did you write it correctly ?

OpenStudy (anonymous):

so its 3 ?

OpenStudy (texaschic101):

is the problem written correctly ?

OpenStudy (anonymous):

yes I rewrote it

OpenStudy (anonymous):

It all depends on the problem written correctly.

OpenStudy (anonymous):

The one you re-wrote still does not work. Can you post an image?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i can try

OpenStudy (texaschic101):

x - y = 33 --> x = 33 + y x - 8y = 9 33 + y - 8y = 9 33 - 7y = 9 33 - 9 = 7y 24 = 7y 24/7 = y 3 3/7 = y

OpenStudy (anonymous):

I think @satellite73 has the correct equations. If you can verify that, then use those equations and try again.

OpenStudy (anonymous):

what we communicate here is failure to have

OpenStudy (anonymous):

@texaschic101 This is a multiple choice question and that is not one of the answers. That is why I think @Kimmieboo1316 has written the problem wrong.

OpenStudy (texaschic101):

Houston we have a problem...

OpenStudy (texaschic101):

I agree

OpenStudy (anonymous):

@satellite73 You are absolutely correct. @Kimmieboo1316 I would love to help you learn this, but the question must be transcribed accurately. If you cannot provide an image, then maybe you can verify the equations provided by @satellite73 .

OpenStudy (anonymous):

I'm trying to provide a picture

OpenStudy (anonymous):

Here's the problem

OpenStudy (anonymous):

My goodness!! No wonder you have such trouble. Please email that image to your teacher so this problem can be properly reported. My suggestion is to go with the equations written by @satellite73 \[x−y=3 \] \[3x−8y=9\]

OpenStudy (anonymous):

Can you solve it using those equations?

OpenStudy (anonymous):

Yes I know and i couldn't its okay though i still passed the homework i got a 70 thank you

OpenStudy (anonymous):

It's not ok because you will build on this topic in later classes. If you want, I can help you learn to solve systems of equations by the substitution method.

OpenStudy (anonymous):

It's okay I will bring it to my teachers attention Monday

OpenStudy (anonymous):

Alright.